what is x in the problem:
28 over 16=x-8 over 20?
i have tried to simplify 28 over 16 to 7 over 4 and then multiply 7 by 5 and subtract 8. when i did this i got 27 over 20, which does not equal to 7 over 4.
i also tried 28 times 20 plus 16 minus 8. this equals 568 which cannot be simplified from 568 over 20 to 7 over 4. Answer by Alan3354(69443) 
 what is x in the problem:
28 over 16=x-8 over 20?
---------------
28/16 = x - 8/20
7/4 = x - 2/5
35/20 = x - 8/20
x = 43/20